# Guide-LLM: An Embodied LLM Agent and Text-Based Topological Map for Robotic Guidance of People with Visual Impairments

### v1

@@ Abstract
Navigation presents a significant challenge for persons with visual impairments (PVI). While traditional aids such as white canes and guide dogs are invaluable, they fall short in delivering detailed spatial information and precise guidance to desired locations. Recent developments in large language models (LLMs) and vision-language models (VLMs) offer new avenues for enhancing assistive navigation. In this paper, we introduce Guide-LLM, an embodied LLM-based agent designed to assist PVI in navigating large indoor environments. Our approach features a novel text-based topological map that enables the LLM to plan global paths using a simplified environmental representation, focusing on straight paths and right-angle turns to facilitate navigation. Additionally, we utilize the LLM’s commonsense reasoning for hazard detection and personalized path planning based on user preferences. Simulated experiments demonstrate the system’s efficacy in guiding PVI, underscoring its potential as a significant advancement in assistive technology. The results highlight Guide-LLM’s ability to offer efficient, adaptive, and personalized navigation assistance, pointing to promising advancements in this field.
@@ I Introduction
Navigating everyday environments can be particularly challenging for persons with visual impairments (PVI), who often depend on specialized tools, help from others, or familiar routes to get around [1, 2, 3, 4, 5]. Traditional aids, such as white canes and guide dogs are essential part of the navigation, however, with the technological developments, further assistance could be possible to improve user confidence in navigation [6]. .
@@
Figure 1: Guide-LLM: Embodied agent consists of text-map, LLM and navigational modules to guide user to destination.
Figure 2: Guide-LLM framework: LLM (Green)Serves as the central controller, using commonsense reasoning to interpret user queries and interact with various modules (Yellow) for decision-making and navigation tasks. Text map (Green) Provides textual representation of the environment used to the path planning module to create route plans. Vector database 1 (Blue) stores static embedding of the environment images, aiding in consistent localization. Vector database 2 (Red) stores navigational image embedding that can be updated or deleted based on the agent’s requirements.
@@
Our framework also incorporates an image retrieval system for localization and a low-level planner to handle robot movement, constraining the robot’s actions to predictable patterns. One of the significant advantages of integrating LLMs into navigation is their ability to leverage commonsense reasoning which has potential for enhancing safety, personalized navigation, and the interpretability of its actions [16, 25], which traditional navigation system often lack. The main contributions of our work are:
  • Introducing a novel embodied LLM agent framework for guiding PVI: We propose Guide-LLM, an innovative framework that utilizes LLMs as embodied agents to assist PVI during navigation.
  • Integration of a novel text-based topological map and image vector database: By combining a text-based topological map with an image vector database, we empower the LLM to carry out high-level planning through commonsense reasoning, minimizing the need for extensive user input or detailed instructions.
  • Simulation evaluation: We validate the effectiveness of our approach through simulations, demonstrating its capability in guiding PVI.
Figure 3: Text map (Left): diagram of text-map, part of the text map is extracted. Example text map representation (Middle): User asks agent to navigate to the elevator. Guide-LLM plans a route(Red line) and begins guiding. Along the route, hazard is detected (wet floor sign), Guide-LLM warns user and suggests alternative path(Green line), Chat-box (Right) shows example communication between Guide-LLM and user
@@ II-A Navigation Aid for People with Visual Impairment
Recent advancements in assistive technology have greatly enhanced the navigation and mobility of PVI. For instance, the robotic white cane [26] employs multimodal sensing and steering assistance to aid users in navigating their surroundings. Additionally, wearable systems have been developed to improve situational awareness by delivering real-time feedback through auditory or haptic signals [27, 28, 29, 30]. [31] introduced a method that allows PVIs to jog on athletic tracks. Despite these advancements, many of these systems are pre-programmed with fixed rules and behaviors, which can limit their adaptability to dynamic scenarios.
To address these limitations, dialogue-based robots have been proposed as an alternative form of assistance, using conversational interfaces to help users reach their destinations [17]. However, these systems still encounter significant challenges, particularly in natural language interaction and adaptability. Their rule-based nature also limits their ability to generalize and adjust to changing environments, which is essential for offering personalized and flexible support to PVI users.
@@ II-B Large Language Models in Robotics
The integration of LLMs into robotics has demonstrated significant potential in areas such as task planning, autonomous driving, multimodal reasoning, and navigation. Frameworks like SayPlan and ReAct enhance robotic capabilities by breaking down complex instructions into actionable sub-tasks and combining reasoning with action [13, 32]. DriveLLM [16] showcased how LLMs can improve interpretability and decision-making in autonomous driving by integrating object-level vector modalities with LLMs, thus enhancing context understanding and explainability in driving scenarios. RoboVQ [33] combined LLMs with vision inputs, enabling robots to perform complex, long-horizon tasks and demonstrating LLMs’ multimodal reasoning capabilities. [34] utilized LLMs as navigation agents, leveraging their reasoning capabilities to develop search heuristics for exploring new environments. SayNav employs LLMs to guide robots through unfamiliar environments by grounding high-level instructions in spatial contexts, highlighting the potential of LLMs in exploration without requiring detailed prior knowledge [35]. [36] demonstrated LLMs’ ability to generalize across tasks, allowing robots to plan and execute complex actions with minimal task-specific training. Additionally, NavGPT [25], MAP-GPT [37], and LGX [38] explore LLM-guided robotic navigation, further extending LLM capabilities in uncharted environments.
Despite these advancements, there is limited research specifically focusing on using LLMs’ commonsense reasoning and contextual understanding to assist people with visual impairments. This gap highlights the need for innovative approaches that harness LLMs’ reasoning abilities to provide more intelligent and adaptive support for PVI.
@@ II-C Vision and Language Navigation
Vision and language navigation (VLN) integrates visual perception with natural language understanding to enable agents to navigate based on spoken or written instructions. A significant challenge in VLN is that natural language instructions often emphasize high-level decisions and landmarks, frequently lacking detailed, low-level movement guidance [39]. Attention-based mechanisms [40, 41, 42] and reinforcement learning approaches [43, 44] have shown promising results in addressing this issue. Recently, the rise of LLMs and VLMs has sparked interest in leveraging these models to enhance VLN capabilities. For instance, LM-Nav [15] uses LLMs to extract landmarks from user queries and VLMs to ground these landmarks in the environment for navigation. Similarly, [45] employ LLMs to translate user queries into actionable navigation tasks and use image segmentation techniques to create a topological map for navigation. [38] combined LLMs commonsense reasoning with VLMs to navigate towards uniquely described objects, showcasing LLMs potential in understanding nuanced language. [46] utilized LLMs to guide the agent by processing action sequences based on text prompts that include navigation instructions, visual landmark descriptions, and the agent’s past trajectory. [17] used LLMs to identify landmarks, describe the environment, and navigate through it.
Despite these advancements, navigation for PVI remains challenging, particularly due to difficulties in describing specific scenes and landmarks, which can impede effective communication with the agent. Our work aims to address this limitation by enabling our agent to infer user intent and automatically resolve ambiguities in their instructions.
@@ III-A1 Agent
We employed GPT-4o [47] as the central component of our framework. GPT-4o is tasked with interpreting user inputs and determining appropriate actions to assist PVI. When a user query is received, the LLM is provided with a system prompt containing instructions to process the query within a specific context. As illustrated in Fig. 4, this prompt is crafted to guide the LLM in assisting the user effectively while addressing potential errors and safety concerns. The system prompt is based on the Chain-of-Thought (CoT) method [48], which enables the LLM to break down tasks into manageable, intermediate steps.
@@ III-A2 Text-Based Topological Map
The framework features a text-based topological map (Fig. 3) that illustrates the spatial relationships between various nodes in the environment (e.g., RoomA is connected to HallwayA, 0.5m north). Nodes are positioned at critical points where the robot needs to turn or take action. This map acts as a schematic reference for the agent, aiding navigation through complex spaces by offering a structured and queryable representation of the environment. It is specifically designed to emphasize straight paths between nodes, as people with visual impairments generally find straight trajectories more intuitive than curved ones [24, 23].
Figure 4: Example of the system prompt.
@@ III-A3 Path Planning Module
This module works in tandem with the agent to plan efficient routes to designated destinations. When the agent initiates a navigation query, the module first consults the text-based topological map to determine potential routes. It calculates the shortest path to the destination and generates a textual description of the route for the agent to process. To enhance flexibility, the module uses a depth-first search (DFS) algorithm to explore all possible routes. This approach allows the agent to assess multiple paths and adapt in real-time based on hazards or user preferences, such as avoiding obstacles, selecting quieter routes, or prioritizing safety concerns. By combining path planning with textual feedback, the agent can make more informed decisions and offer a personalized navigation experience for the user.
@@ III-A4 Embedding Module
We use pre-trained CLIP [10] to generate embeddings from visual data, which are then stored in a vector database alongside relevant metadata, including location and orientation information. This metadata allows the LLM to reference the location and orientation of specific images upon retrieval. The embedding module is essential for localization and subgoal selection, as it supports both text-to-image and image-to-image similarity searches. This functionality enables the agent to localize itself accurately and select appropriate subgoals.
@@ III-A5 Vector Database
The vector database manages high-dimensional vector representations of the environment, which are generated by the embedding module. This setup enables efficient comparison of current observations with previously stored data. The database is crucial for real-time decision-making, allowing the LLM to retrieve images that match current observations and navigation-related images. Two separate vector databases are maintained: one for storing all environment images and another for images related to ongoing navigation. This separation helps minimize ambiguity and ensures that navigation-related images are retrieved more accurately, preventing confusion from unrelated images in a single, densely populated database.
@@ III-A6 Low-Level Planner
The low-level planner converts the high-level decisions made by the agent into executable actions for the robot. It ensures that these commands are physically feasible, considering the robot’s movement capabilities. By connecting high-level decision-making with practical execution, the low-level planner enables the robot to operate safely and efficiently within various environments.
@@ III-B Global Path Planning
Figure 2 illustrates the overall process of the proposed framework. We start with the assumption that the text map and images in the vector database are pre-labeled. The process begins with the LLM processing the user’s query and the system prompt. The LLM then generates high-level plans, image query commands, and responses for the user.
As shown in Fig. 2, each output is communicated separately for clear communication between modules. High-level plans and user responses are delivered through a voice-to-text interface, while image queries retrieve navigation-related images from the vector database. These images are then embedded into a secondary vector database, which refines the navigation process by reducing ambiguity and ensuring that only relevant images are retrieved.
@@ III-C Topological Navigation
The agent begins navigation by querying the vector database to retrieve the image for the next node. This image is compared to the current observation using cosine similarity for localization. To improve place recognition accuracy, a similarity check is performed each time the robot has traveled a specified distance or made a turn. The distance to the next node is measured by the robot’s odometer through the low-level planner, and a message is sent to the agent upon arrival. If the similarity score exceeds a predefined threshold, the LLM concludes that the target node has been reached and generates the next set of movement commands to proceed. This process continues until the final destination is reached.
@@ III-D Localization and Error Handling
Localization in our agent framework employs a dual-layered approach. Initially, the agent verifies its arrival at the desired node by querying the navigational vector database and comparing the retrieved image with the robot’s current observation. If the similarity score between these images falls below a set threshold, the agent detects a potential localization error, suggesting that the robot may be in the wrong location. To address this, the agent initiates a broader search by querying the main vector database, which contains embeddings of the entire environment. This fallback mechanism allows the system to re-localize accurately by finding the most relevant match among all environment images. This dynamic error-handling process ensures that even if the robot is initially misaligned or disoriented, the system can correct its location in real-time, maintaining the desired path for safe and efficient navigation.
@@ III-F Personalization Potential
A key strength of our agent is its ability to personalize the navigation experience according to each user’s specific preferences and needs. The agent’s natural language interaction capabilities enable it to adjust its behavior in real-time, facilitating this personalization. For example, it can modify its path-planning to match a user’s preferred walking speed, route preferences (e.g., avoiding stairs or choosing quieter areas), or specific safety concerns, such as steering clear of potential hazards. Additionally, the system can incorporate feedback from previous interactions, progressively refining its decisions to align with the user’s habits and preferences. For instance, if a user consistently opts for longer, less crowded routes over shorter ones, the agent can integrate this preference into its future planning. This adaptability extends beyond navigation; the agent can engage in personalized dialogue, adjusting the level of detail and communication style to meet different users’ needs, whether they prefer brief instructions or more detailed explanations.
@@ IV Results
We evaluated the capabilities of our Guide-LLM by testing its ability to guide visually impaired individuals through simulated environments. The simulations are divided into four key scenarios, each designed to highlight distinct aspects of the LLM’s utility in navigation, decision-making, error handling, and hazard detection. All tests were conducted in a simulated environment using the iGibson simulator with a TurtleBot platform.
TABLE I: Success Rates of Guide-LLM Navigation with Different Configurations
Environment All components No System Prompt No Path Planning Module
Large House 90% 0% 33.33%
Office 83.33% 0% 40%
TABLE II: success rate of Localization error detection and recovery
Error Scenario Success Rate %

### v2

@@ Abstract
Navigation presents a significant challenge for persons with visual impairments (PVI). While traditional aids such as white canes and guide dogs are invaluable, they fall short in delivering detailed spatial information and precise guidance to desired locations. Recent developments in large language models (LLMs) and vision-language models (VLMs) offer new avenues for enhancing assistive navigation. In this paper, we introduce Guide-LLM, an embodied LLM-based agent designed to assist PVI in navigating large indoor environments. Our approach features a novel text-based topological map that enables the LLM to plan global paths using a simplified environmental representation, focusing on straight paths and right-angle turns to facilitate navigation. Additionally, we utilize the LLM’s commonsense reasoning for hazard detection and personalized path planning based on user preferences. Simulated experiments demonstrate the system’s efficacy in navigating indoor environments based on user queries, underscoring its potential as a significant advancement in assistive technology. The results highlight Guide-LLM’s ability to offer efficient, adaptive, and personalized navigation assistance, pointing to promising advancements in this field.
@@
Code and replication details: https://github.com/awd1779/Guide-LLM
@@ I Introduction
Navigating everyday environments can be particularly challenging for persons with visual impairments (PVI), who often depend on specialized tools, help from others, or familiar routes to get around [1, 2, 3, 4, 5]. Traditional aids, such as white canes and guide dogs, are essential parts of the navigation; however, with the technological developments, further assistance could be possible to improve user confidence in navigation [6].
@@
Figure 1: Guide-LLM: The embodied agent consists of a text map, LLM, and navigational modules to guide the user to the destination.
Figure 2: Guide-LLM framework: LLM (green) serves as the central controller, using commonsense reasoning to interpret user queries and interact with various modules (yellow) for decision-making and navigation tasks. Text map (Green) provides a textual representation of the environment used by the path planning module to create route plans. Vector database 1 (blue) stores static embeddings of the environment images, aiding in consistent localization. Vector database 2 (red) stores navigational image embedding that can be updated or deleted based on the agent’s requirements.
@@
Guide-LLM also incorporates an image retrieval system for localization and a low-level planner to handle robot movement, constraining the robot’s actions to predictable patterns. One of the significant advantages of integrating LLMs into navigation is their ability to leverage commonsense reasoning, which has the potential for enhancing safety, personalized navigation, and the interpretability of its actions [16, 25], which traditional navigation systems often lack. The main contributions of our work are:
  • We propose Guide-LLM, an innovative framework that employs LLM as embodied agents to assist visually impaired individuals in navigation.
  • By integrating a novel text-based topological map, LLM, and navigational tools (Fig. 1), our approach enables high-level planning while minimizing the need for users to provide extensive descriptions of destination.
  • In comprehensive simulations, we validate our method’s localization error detection and recovery, personalization potential, and hazard detection demonstrating the robust capabilities required for guiding PVI.
Figure 3: Text map (Left): diagram of text map, part of the text map is extracted. Example text map representation (Middle): User asks the agent to navigate to the elevator. Guide-LLM plans a route (red line) and begins guiding. Along the route, a hazard is detected (wet floor sign), Guide-LLM warns the user and suggests an alternative path (green line), and the chat box (right) shows an example communication between Guide-LLM and the user.
@@ II-A Navigation Aid for People with Visual Impairment
Recent advancements in assistive technology have greatly enhanced the navigation and mobility of PVI. For instance, the robotic white cane [26] employs multimodal sensing and steering assistance to aid users in navigating their surroundings. Additionally, wearable systems have been developed to improve situational awareness by delivering real-time feedback through auditory or haptic signals [27, 28, 29, 30]. [31] introduced a method that allows PVIs to jog on athletic tracks. Despite these advancements, many of these systems are pre-programmed with fixed rules and behaviors, which can limit their adaptability to dynamic scenarios. To address these limitations, dialogue-based robots have been proposed as an alternative form of assistance, using conversational interfaces to help users reach their destinations [17]. However, these systems still encounter significant challenges, particularly in natural language interaction and adaptability. Their rule-based nature also limits their ability to generalize and adjust to changing environments, which is essential for offering personalized and flexible support to PVI users.
@@ II-B Large Language Models in Robotics
The integration of LLMs into robotics has demonstrated significant potential in areas such as task planning, autonomous driving, multimodal reasoning, and navigation. Frameworks like SayPlan and ReAct enhance robotic capabilities by breaking down complex instructions into actionable sub-tasks and combining reasoning with action [13, 32]. DriveLLM [16] showcased how LLMs can improve interpretability and decision-making in autonomous driving by integrating object-level vector modalities with LLMs, thus enhancing context understanding and explainability in driving scenarios. RoboVQ [33] combined LLMs with vision inputs, enabling robots to perform complex, long-horizon tasks and demonstrating LLMs’ multimodal reasoning capabilities. [34] utilized LLMs as navigation agents, leveraging their reasoning capabilities to develop search heuristics for exploring new environments. SayNav employs LLMs to guide robots through unfamiliar environments by grounding high-level instructions in spatial contexts, highlighting the potential of LLMs in exploration without requiring detailed prior knowledge [35]. [36] demonstrated LLMs’ ability to generalize across tasks, allowing robots to plan and execute complex actions with minimal task-specific training. Additionally, NavGPT [25], MAP-GPT [37], and LGX [38] explore LLM-guided robotic navigation, further extending LLM capabilities in uncharted environments. Despite these advancements, there is limited research specifically focusing on using LLMs’ commonsense reasoning and contextual understanding to assist people with visual impairments. This gap highlights the need for innovative approaches that harness LLMs’ reasoning abilities to provide more intelligent and adaptive support for PVI.
@@ II-C Vision and Language Navigation
Vision and language navigation (VLN) integrates visual perception with natural language understanding to enable agents to navigate based on spoken or written instructions. A significant challenge in VLN is that natural language instructions often emphasize high-level decisions and landmarks, frequently lacking detailed, low-level movement guidance [39]. Attention-based mechanisms [40, 41, 42] and reinforcement learning approaches [43, 44] have shown promising results in addressing this issue. Recently, the rise of LLMs and VLMs has sparked interest in leveraging these models to enhance VLN capabilities. For instance, LM-Nav [15] uses LLMs to extract landmarks from user queries and VLMs to ground these landmarks in the environment for navigation. Similarly, [45] employs LLMs to translate user queries into actionable navigation tasks and use image segmentation techniques to create a topological map for navigation. [38] combined LLMs commonsense reasoning with VLMs to navigate towards uniquely described objects, showcasing LLMs potential in understanding nuanced language. [46] utilized LLMs to guide the agent by processing action sequences based on text prompts that include navigation instructions, visual landmark descriptions, and the agent’s past trajectory. [17] used LLMs to identify landmarks, describe the environment, and navigate through it. Despite these advancements, navigation for PVI remains challenging, particularly due to difficulties in describing specific scenes and landmarks, which can impede effective communication with the agent. Our work aims to address this limitation by enabling our agent to infer user intent and automatically resolve ambiguities in their instructions.
TABLE I: Feature Comparison Table of Our Method and Others
Approach Natural language interactions Doesn’t Require Detailed Goal Knowledge Localization Error & Recovery Identify potential Hazards Personalization
Guide-LLM
(Our Approach)
Traditional Navigation Aid
[26][47]
Wearable System
[27, 28, 29, 30]
VLN Approach
[15][17][38][45]
@@ III-A1 Agent
We used GPT-4o [48] as the central component of our framework. As shown in Fig. 2, it is tasked with interpreting user inputs and determining appropriate actions to assist the user. When a user query is received, the LLM is provided with a system prompt containing instructions to process the query within a specific context. The system prompt designed to provide context to LLM which which enables the LLM to perform appropriate action in each steps.
@@ III-A2 Text-Based Topological Map
To generate our text-based topological map (Fig. 3), we first construct an occupancy grid map of the area and then manually place nodes at each intersection, turn-able location, and critical navigation decision point. These nodes form the vertex set $V$ of a directed and weighted graph $\mathcal{G}=(V,E)$, where each edge $e_{ij}$ connects node $i$ to node $j$ and is annotated with a distance $d_{ij}$ and a relative direction $\theta_{ij}$. The straight edges that connect these nodes not only reflect direct traversable paths but also encapsulate the metric and directional information necessary for navigation. When embedded in a Euclidean plane, each node is assigned a coordinate $p_{i}=(x_{i},y_{i})$, and the spatial transformation from node $i$ to node $j$ is expressed by
$p_{j}=p_{i}+\left(d_{ij}\cos\theta_{ij},\;d_{ij}\sin\theta_{ij}\right).$
We intentionally avoid curved paths or minor direction changes to maintain simplicity, in accordance with findings [24, 23] that suggest straight paths are more comfortable for individuals with visual impairments.
@@ III-A3 Path Planning Module
The planning module works in tandem with the agent to plan efficient routes to designated destinations. Upon a navigation query, it processes the text-based topological map, and generates a textual description of the route for the agent to process. This approach allows the agent to assess multiple paths and adapt in real-time based on hazards or user preferences.
@@ III-A4 Embedding Module
We use off-the-shelf CLIP [10] to generate embeddings from visual data, which are then stored in a vector database alongside location and orientation metadata. This metadata allows the LLM to reference the location and orientation of specific images upon retrieval. The embedding module is essential for localization and sub-goal selection, as it supports both text-to-image and image-to-image similarity searches.
@@ III-A5 Vector Database
The vector database stores high-dimensional representations of the environment generated by the embedding module, enabling real-time comparison of current observations with previously stored data. The agent accesses this data by querying navigation-related images and uses them to compare current observations with retrieved images.
@@ III-A6 Low-Level Planner
The low-level planner converts the high-level decisions made by the agent into executable actions for the robot. It ensures that these commands are physically feasible, considering the robot’s movement capabilities.
@@ III-B Global Path Planning
Figure 2 illustrates the overall process of the proposed framework. We start with the assumption that the text map and images in the main vector database are pre-labeled. The process begins with the agent processing the user’s query and the system prompt. The agent then generates high-level plans, navigational image query commands, and responses for the user. As shown in Fig. 2, each output is communicated separately for clear communication between modules. High-level plans and user responses are delivered through a voice-to-text interface, while image queries retrieve navigation-related images from the vector database. These images are then embedded into a secondary vector database, which refines the navigation process by reducing ambiguity and ensuring that only relevant images are retrieved.
@@ III-C Topological Navigation
The agent begins navigation by querying the secondary vector database to retrieve the image corresponding to the current node. The retrieved image is compared to the current observation using cosine similarity. To improve place recognition, a similarity check is performed each time the robot travels to the next node or makes a turn. The low-level planner measures the distance traveled to the next node, and upon arrival, it sends a message to the agent. If the similarity score exceeds a predefined threshold, the agent concludes that the target node has been reached and generates the next set of movement commands. This process continues until the final destination is reached.
@@ III-D Localization and Error Handling
Localization in the Guide-LLM framework employs a dual-layered approach. Initially, the agent verifies its arrival at the desired node by querying the navigational vector database and comparing the retrieved image with the robot’s current observation. If the similarity score between these images falls below a set threshold, the agent detects a potential localization error, suggesting that the robot may be in the wrong location. To address this, the agent initiates a broader search by querying the main vector database, which contains embeddings of the entire environment. This fallback mechanism allows Guide-LLM to re-localize by finding the most relevant match among all environment images. This dynamic error-handling process ensures that even if an agent is initially misaligned or disoriented, Guide-LLM can correct its location in real-time, maintaining the desired path for navigation.
@@ III-F Personalization Potential
A key strength of Guide-LLM is its ability to personalize the navigation experience according to each user’s specific preferences and needs. The agent’s natural language interaction capabilities enable it to adjust its behavior in real-time, facilitating this personalization. For example, it can modify its path-planning to match a user’s preferred walking speed, route preferences (e.g., avoiding stairs or choosing quieter areas), or specific safety concerns, such as steering clear of potential hazards. Additionally, Guide-LLM can incorporate feedback from previous interactions, progressively refining its decisions to align with the user’s habits and preferences. For instance, if a user consistently opts for longer, less crowded routes over shorter ones, the agent can integrate this preference into its future planning. This adaptability extends beyond navigation; the agent can engage in personalized dialogue, adjusting the level of detail and communication style to meet different users’ needs, whether they prefer brief instructions or more detailed explanations.
TABLE II: LLM Performance Comparison on Navigation (Success Rates)
Model (Parameters) House Office
GPT-4o [48] 84% 86%
Claude 3.5 Sonnet [49] 74% 78%
Llama 3.2 Vision 90B [50] 76% 74%
Llama 3.2 Vision 11B [50] 20% 24%
Qwen2.5-VL-7B [51] 0% 4%
@@ IV Results
We evaluated the Guide-LLM by testing its ability to navigate through a simulated indoor environment based on user queries. The simulations are divided into scenarios designed to highlight distinct aspects of the Guide-LLM’s utility in navigation, decision-making, error handling, and hazard detection and personalization.
IV-A Experiment Setup
We used iGibson [52] simulator to validate Guide-LLM. The simulator was deployed using AWS G4dn.2xlarge instance and LLM models were deployed in the cloud using API.
IV-B Model Selection and Topological Navigation
To assess different LLM’s ability to reach specified destinations and select optimal LLM for Guide-LLM, we performed evaluations in two indoor environments within the iGibson simulator [49]. These environments included a three-room house and a larger office-like setting with multiple hallways and offices, as shown in Fig. 3. In both environments, the agent needed to reach a specified destination using four key components: (1) a user query, (2) a text-based topological map, (3) a vector database, and (4) its own reasoning ability. To examine the impact of model scale on navigation performance, we selected five vision-enabled LLMs: three models with over 90 billion parameters and two models with fewer than 11 billion parameters. We repeated each trial 50 times per environment to ensure statistical robustness. If the agent successfully arrived at the destination, we continued issuing new navigation queries until it eventually failed to reach a subsequent destination. If the agent failed to reach its current destination, we reset the robot’s position and counted that attempt as a failure. Table II shows that GPT-4o[48] achieved the highest overall success rate. Most navigation failures resulted from the LLM “forgetting” which command it should generate or confused with turning directions, suggesting that limited real-world grounding can hinder consistent decision-making. Furthermore, our experiments reveal that model size plays a crucial role in navigation success rates: GPT-4o’s performance was approximately three times higher than that of Llama 3.2 Vision 11B, suggesting a positive correlation between parameter count and navigation capability.
TABLE III: Localization Error Detection and Recovery
Error Scenario Success Rate (%)